# MME: A Comprehensive Evaluation Benchmark for Multimodal Large Language Models

### v4

@@ Abstract
Multimodal Large Language Model (MLLM) relies on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in recent studies, such as writing poems based on an image. However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation. In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME [^1] MME is collected by Xiamen University for academic research only. This is the v1 version and will be updated as MLLMs evolve.. It measures both perception and cognition abilities on a total of 14 subtasks. In order to avoid data leakage that may arise from direct use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed. The concise instruction design allows us to fairly compare MLLMs, instead of struggling in prompt engineering. Besides, with such an instruction, we can also easily carry out quantitative statistics. A total of 30 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization. The data application manner and online leaderboards are released at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
@@ 1 Introduction
The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) [37, 8, 20, 23, 13, 9]. It refers to using LLM as a brain to process multimodal information and give reasoning results [49]. Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks [13, 45, 38, 27, 54, 31]. The three representative abilities of LLM [55], including instruction following [40], In-Context Learning (ICL) [10], and Chain-of-Thought (CoT) [43] are also manifested in multimodality. For example, Flamingo [8] turns on multimodal ICL, which can adapt to new tasks by giving a few examples. PaLM-E [13] achieves amazing OCR-free math reasoning via CoT. GPT-4V [37] shows even more ability in a variety of complex reasoning tasks [46]. MiniGPT-4 [59] implements GPT-4[37]-like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning. These emergent abilities of MLLMs are exciting and imply that a new dawn has broken in artificial intelligence.
@@
Figure 2: Leaderboards on our MME benchmark. (1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800. (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200. The score is the sum of the accuracy and the accuracy+ in Tables. 1 and 2. A total of 30 advanced MLLMs joint the leaderboards. For the sake of presentation, we only show 10 models for each list, in which the top three ones are given clear trophy logos.
@@
Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects. The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance. Specifically, the first manner [47, 12, 42] evaluates on existing traditional multimodal datasets, such as image caption [11] and VQA [16, 36, 32]. However, on the one hand, it may be hard to reflect the emergent abilities of MLLMs on these datasets. On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training. The second manner [48] is to collect data for an open-ended evaluation, but either the data is unavailable to public by now [57] or the amount is small (only 50 images) [48]. The third manner focuses on one aspect of MLLMs, such as object hallucination [25] or adversarial robustness [56], which is powerless to comprehensive evaluation.
@@
In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs. We argue that a universal comprehensive evaluation benchmark should have the following four characteristics: (1) It should cover as much as possible, including both perception and cognition abilities. The former refers to recognizing the specific object, such as its existence, count, position, and color. The latter refers to compositing the perception information and the knowledge in LLM to deduce more complex answers. It is obvious that the former is the premise of the latter. (2) Its data or annotations should not come from existing publicly available datasets as much as possible, avoiding the risk of data leakage. (3) Its instructions should be as concise as possible and in line with human cognition. Although instruction design may have a large impact on the output, all models should be tested under the same unified instructions for fair comparison. A good MLLM should be able to generalize to such concise instructions. (4) The responses of MLLMs to the instructions should be intuitive and convenient for quantitative analysis. The open-ended answer of MLLMs poses significant challenges to the quantization. Existing methods tend to use GPT or manual scoring [21, 29, 48], but there may be problems of inaccuracy and subjectivity.
@@
We conduct massive experiments to evaluate the zero-shot performance of 30 advanced MLLMs on the 14 subtasks. The evaluated MLLMs include BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6]. As displayed in Fig. 2 that consists of 2 overall leaderboards (perception and cognition) and 14 individual leaderboards, these MLLMs show clear discrepancies in our MME evaluation benchmark. Fig. 3 also provides a comparison from the other perspective. We can see the range that current MLLMs can reach in each capability dimension. More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination [50, 25], as shown in Fig. 4. It is expected that these findings are instructive for the subsequent model optimization.
@@ 2.3.1 Perception Tasks
We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem [50, 25]. That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig. 4.
Coarse-Grained Recognition. The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position. The images are sampled from COCO [26], but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations. Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets. This requires MLLMs to be able to understand the instructions and infer corresponding answers. In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
Fine-Grained Recognition. The fine-grained recognition is more about testing the knowledge resources of MLLMs. The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively. For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]? Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets [19, 34, 35, 58, 44] and all of the instructions are manually designed.
OCR. Optical Character Recognition (OCR) is also a foundational capability of MLLMs, serving for subsequent text-based tasks such as text translation and text understanding. The images are sampled from [30] and all of the instruction-answer pairs are manually designed. Considering that MLLMs are still in its infancy, we only choose the relatively simple samples in this version of MME. The numbers of image and instruction-answer pairs are 20 and 40, respectively.
@@ 2.3.2 Cognition Tasks
Commonsense Reasoning. Unlike the ScienceQA dataset [32] that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life. For example, given a photo of a down jacket, asking MLLMs whether it is appropriate to wear the cloth when it is cold (or hot). These are basic knowledge that humans can judge instantly without complex step-by-step reasoning. Therefore, we expect MLLMs to perform well in a zero-short setting. The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed. There are a total of 70 images and 140 instruction-answer pairs.
Numerical Calculation. It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in [20]. In this version, we only consider relatively easy arithmetic problems, such as addition and multiplication. There are 20 images and 40 instruction-answer pairs. The images are all manually taken, and the instruction-answer pairs are all manually designed.
Text Translation. Considering that the MLLM [5] supports both English and Chinese, we set the text translation subtask. It requires MLLMs to translate the Chinese written in an image to the corresponding English. In this version, we only design basic translation problems, which will be updated according to the development of MLLMs in the future. The images of this part are all manually taken, and the instruction-answer pairs are all manually designed. There are a total of 20 images and 40 instruction-answer pairs.
Code Reasoning. It requires MLLMs to read the code in the images and automatically complete logical operation inside the code. A similar task that writes website code based on an image has been demonstrated in [59]. The images are all manually taken, and the instruction-answer pairs are all manually designed. We only set basic code problems in this version. There are in total 20 images and 40 instruction-answer pairs.
@@ 3 Experiments
In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 [23], InstructBLIP [12], MiniGPT-4 [59], PandaGPT [39], Multimodal-GPT [15], VisualGLM-6B [5], ImageBind-LLM [17], VPGTrans [53], LaVIN [33], mPLUG-Owl [48], Octopus [3], Muffin [51], Otter [22], LRV-Instruction [28], Cheetor [24], LLaMA-Adapter-v2 [14], GIT2 [41], BLIVA [18], Lynx [52], MMICL [54], GPT-4V [37], Skywork-MM [4], mPLUG-Owl2 [48], Qwen-VL-Chat [9], XComposer-VL [7], LLaVA [29], Lion [2], SPHINX [27], InfMLLM [1], and WeMM [6].
@@ 4 Analysis
The third problem is a lack of reasoning. In the third row of Fig. 4, we can see from the red text that the MLLM already knows that the first image is not an office place, but still gives an incorrect answer of “yes”. Analogously, in the second image, the MLLM has calculated the right arithmetic result, but finally delivers a wrong answer. These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs. Adding CoT prompts, such as “Let’s think step by step” [13], may yield better results. We look forward to a further in-depth research.
@@ 5 Conclusion
This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics. 30 advanced MLLMs are evaluated on MME and the experimental results show that there is still a large room to improve. We also summarize the common problem raised in experimental results, providing valuable guidance for the development of MLLM.

### v5

@@ Abstract
Multimodal Large Language Model (MLLM) relies on the powerful LLM to perform multimodal tasks, showing amazing emergent abilities in recent studies, such as writing poems based on an image. However, it is difficult for these case studies to fully reflect the performance of MLLM, lacking a comprehensive evaluation. In this paper, we fill in this blank, presenting the first comprehensive MLLM Evaluation benchmark MME. It measures both perception and cognition abilities on a total of 14 subtasks. In order to avoid data leakage that may arise from direct use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed. The concise instruction design allows us to fairly compare MLLMs, instead of struggling in prompt engineering. Besides, with such an instruction, we can also easily carry out quantitative statistics. A total of 30 advanced MLLMs are comprehensively evaluated on our MME, which not only suggests that existing MLLMs still have a large room for improvement, but also reveals the potential directions for the subsequent model optimization. The data are released at the project page: https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
@@ 1 Introduction
The thriving of Large Language Model (LLM) has paved a new road to the multimodal field, i.e., Multimodal Large Language Model (MLLM) openai2023gpt4 ; huang2023language ; li2023blip ; driess2023palm ; team2025kimi ; tu2024overview . It refers to using LLM as a brain to process multimodal information and give reasoning results yin2023survey . Equipped with the powerful LLM, MLLM is expected to address more complex multi-modal tasks driess2023palm ; wen2025dycrowd ; shen2023hugginggpt ; lin2023sphinx ; zhang2025speechact ; zhao2023mmicl ; liu2023mmbench ; liu2021geometry ; zhang2025logavatar ; ying2024mmt ; fang2024mmbench .The three representative abilities of LLM zhao2023survey , including instruction following touvron2023llama , In-Context Learning (ICL) brown2020language , and Chain-of-Thought (CoT) wei2022chain are also manifested in multimodality. For example, Flamingo alayrac2022flamingo turns on multimodal ICL, which can adapt to new tasks by giving a few examples. PaLM-E driess2023palm achieves amazing OCR-free math reasoning via CoT. GPT-4V openai2023gpt4 shows even more ability in a variety of complex reasoning tasks wu2023early . MiniGPT-4 zhu2023minigpt implements GPT-4openai2023gpt4 -like instruction following capabilities, such as converting images into corresponding website codes, by introducing multimodal instruction tuning. These emergent abilities of MLLMs are exciting and imply that a new dawn has broken in artificial intelligence.
@@
Figure 2: Leaderboards on our MME benchmark. (1) and (2) are the overall leaderboards of perception and cognition respectively, in which the full score of the former is 2000 and that of the latter is 800. (3)-(16) are the leaderboards of the 14 subtasks with the full score of 200. The score is the sum of the accuracy and the accuracy+ in Tables 1 and 2. A total of 30 advanced MLLMs joint the leaderboards. For the sake of presentation, we only show 10 models for each list, in which the top three ones are given clear trophy logos.
@@
Although these models exhibit surprising conversational capabilities when conducting everyday chats, we still know little about how well they quantitatively perform in various aspects. The existing three common quantitative evaluation manners for MLLMs have their limitations that are difficult to comprehensively evaluate performance. Specifically, the first manner xu2022multiinstruct ; dai2023instructblip ; wang2023visionllm evaluates on existing traditional multimodal datasets, such as image caption chen2015microsoft and VQA goyal2017making ; marino2019ok ; lu2022learn . However, on the one hand, it may be hard to reflect the emergent abilities of MLLMs on these datasets. On the other hand, since the training sets of large models are no longer unified, it is difficult to guarantee that all MLLMs have not used the testing set for training. The second manner ye2023mplug is to collect data for an open-ended evaluation, but either the data is unavailable to public by now zhao2023chatbridge or the amount is small (only 50 images) ye2023mplug . The third manner focuses on one aspect of MLLMs, such as object hallucination li2023evaluating or adversarial robustness zhao2023evaluating , which is powerless to comprehensive evaluation.
@@
In light of these concerns, a new comprehensive evaluation benchmark is urgently needed to match the flourish of MLLMs. We argue that a universal comprehensive evaluation benchmark should have the following four characteristics: (1) It should cover as much as possible, including both perception and cognition abilities. The former refers to recognizing the specific object, such as its existence, count, position, and color. The latter refers to compositing the perception information and the knowledge in LLM to deduce more complex answers. It is obvious that the former is the premise of the latter. (2) Its data or annotations should not come from existing publicly available datasets as much as possible, avoiding the risk of data leakage. (3) Its instructions should be as concise as possible and in line with human cognition. Although instruction design may have a large impact on the output, all models should be tested under the same unified instructions for fair comparison. A good MLLM should be able to generalize to such concise instructions. (4) The responses of MLLMs to the instructions should be intuitive and convenient for quantitative analysis. The open-ended answer of MLLMs poses significant challenges to the quantization. Existing methods tend to use GPT or manual scoring li2023mimic ; liu2023visual ; ye2023mplug , but there may be problems of inaccuracy and subjectivity.
@@
We conduct massive experiments to evaluate the zero-shot performance of 30 advanced MLLMs on the 14 subtasks. The evaluated MLLMs include BLIP-2 li2023blip , InstructBLIP dai2023instructblip , MiniGPT-4 zhu2023minigpt , PandaGPT su2023pandagpt , Multimodal-GPT gong2023multimodal , VisualGLM-6B git-visualglm-6b , ImageBind-LLM han2023imagebind , VPGTrans zhang2023transfer , LaVIN luo2023cheap , mPLUG-Owl ye2023mplug , Octopus git-octopus , Muffin yu2023reformulating , Otter li2023otter , LRV-Instruction liu2023aligning , Cheetor li2023finetuning , LLaMA-Adapter-v2 gao2023llama , GIT2 wang2022git , BLIVA hu2023bliva , Lynx zeng2023matters , MMICL zhao2023mmicl , GPT-4V openai2023gpt4 , Skywork-MM git-skywork , mPLUG-Owl2 ye2023mplug , Qwen-VL-Chat bai2023qwen , XComposer-VL git-xcomposer , LLaVA liu2023visual , Lion git-lion , SPHINX lin2023sphinx , InfMLLM git-infmllm , and WeMM git-wemm . As displayed in Fig. 2 that consists of 2 overall leaderboards (perception and cognition) and 14 individual leaderboards, these MLLMs show clear discrepancies in our MME evaluation benchmark. Fig. 3 also provides a comparison from the other perspective. We can see the range that current MLLMs can reach in each capability dimension. More importantly, we have summarized four prominent problems exposed in experiments, including inability to follow basic instructions, a lack of basic perception and reasoning, as well as object hallucination yin2023woodpecker ; li2023evaluating , as shown in Fig. 4. It is expected that these findings are instructive for the subsequent model optimization.
@@ 2.3.1 Perception Tasks
We argue that perception is one of the most fundamental capabilities of MLLMs, and the lack of perception will easily lead to the object hallucination problem yin2023woodpecker ; li2023evaluating . That is, MLLM will answer questions based on its own fantasies rather than based on the realistic content of the image, as displayed in Fig. 4.
Coarse-Grained Recognition. The contents of coarse-grained recognition include the existence of common objects, and their count, color, and position. The images are sampled from COCO lin2014microsoft , but the instruction-answer pairs are all manually constructed, rather than directly using publicly available annotations. Even if MLLMs have seen these COCO images, our manually prepared pairs are not presented in their training sets. This requires MLLMs to be able to understand the instructions and infer corresponding answers. In each perception subtask of existence, count, color, and position, we prepare 30 images with 60 instruction-answer pairs.
Fine-Grained Recognition. The fine-grained recognition is more about testing the knowledge resources of MLLMs. The subtasks consist of recognizing movie posters, celebrities, scenes, landmarks, and artworks, containing 147, 170, 200, 200, and 200 images respectively. For the celebrities, we plot a red box to a person with a clearly visible face in the image, and the corresponding instruction is “Is the actor inside the red box named [celebrity name]? Please answer yes or no.” Similar with the above coarse-grained recognition, the images of these subtasks are from publicly available datasets huang2020movienet ; mao2017deepart ; mao2019visual ; zhou2014learning ; weyand2020google and all of the instructions are manually designed.
OCR. Optical Character Recognition (OCR) is also a foundational capability of MLLMs, serving for subsequent text-based tasks such as text translation and text understanding. The images are sampled from liu2019curved and all of the instruction-answer pairs are manually designed. Considering that MLLMs are still in its infancy, we only choose the relatively simple samples in this version of MME. The numbers of image and instruction-answer pairs are 20 and 40, respectively.
@@ 2.3.2 Cognition Tasks
Commonsense Reasoning. Unlike the ScienceQA dataset lu2022learn that requires specialized knowledge, the commonsense refers to the basic knowledge in daily life. For example, given a photo of a down jacket, asking MLLMs whether it is appropriate to wear the cloth when it is cold (or hot). These are basic knowledge that humans can judge instantly without complex step-by-step reasoning. Therefore, we expect MLLMs to perform well in a zero-short setting. The images are all manually photographed or generated by diffusion models, and the instruction-answer pairs are all manually designed. There are a total of 70 images and 140 instruction-answer pairs.
Numerical Calculation. It requires MLLMs to be able to read the arithmetic problem in the image and output the answer in an end to end way, which has been demonstrated in huang2023language . In this version, we only consider relatively easy arithmetic problems, such as addition and multiplication. There are 20 images and 40 instruction-answer pairs. The images are all manually taken, and the instruction-answer pairs are all manually designed.
Text Translation. Considering that the MLLM git-visualglm-6b supports both English and Chinese, we set the text translation subtask. It requires MLLMs to translate the Chinese written in an image to the corresponding English. In this version, we only design basic translation problems, which will be updated according to the development of MLLMs in the future. The images of this part are all manually taken, and the instruction-answer pairs are all manually designed. There are a total of 20 images and 40 instruction-answer pairs.
Code Reasoning. It requires MLLMs to read the code in the images and automatically complete logical operation inside the code. A similar task that writes website code based on an image has been demonstrated in zhu2023minigpt . The images are all manually taken, and the instruction-answer pairs are all manually designed. We only set basic code problems in this version. There are in total 20 images and 40 instruction-answer pairs.
@@ 3 Experiments
In this section, a total of 30 MLLMs are evaluated on our MME benchmark, including BLIP-2 li2023blip , InstructBLIP dai2023instructblip , MiniGPT-4 zhu2023minigpt , PandaGPT su2023pandagpt , Multimodal-GPT gong2023multimodal , VisualGLM-6B git-visualglm-6b , ImageBind-LLM han2023imagebind , VPGTrans zhang2023transfer , LaVIN luo2023cheap , mPLUG-Owl ye2023mplug , Octopus git-octopus , Muffin yu2023reformulating , Otter li2023otter , LRV-Instruction liu2023aligning , Cheetor li2023finetuning , LLaMA-Adapter-v2 gao2023llama , GIT2 wang2022git , BLIVA hu2023bliva , Lynx zeng2023matters , MMICL zhao2023mmicl , GPT-4V openai2023gpt4 , Skywork-MM git-skywork , mPLUG-Owl2 ye2023mplug , Qwen-VL-Chat bai2023qwen , XComposer-VL git-xcomposer , LLaVA liu2023visual , Lion git-lion , SPHINX lin2023sphinx , InfMLLM git-infmllm , and WeMM git-wemm .
@@ 4 Analysis
The third problem is a lack of reasoning. In the third row of Fig. 4, we can see from the red text that the MLLM already knows that the first image is not an office place, but still gives an incorrect answer of “yes”. Analogously, in the second image, the MLLM has calculated the right arithmetic result, but finally delivers a wrong answer. These phenomena indicate that the logic chain is broken during the reasoning process of MLLMs. Adding CoT prompts, such as “Let’s think step by step” driess2023palm , may yield better results. We look forward to a further in-depth research.
@@ 5 Conclusion
This paper has presented the first MLLM evaluation benchmark MME that has four distinct characteristics in terms of task type, data source, instruction design, quantitative statistics. 30 advanced MLLMs are evaluated on MME and the experimental results show that there is still a large room to improve. We also summarize the common problem raised in experimental results, providing valuable guidance. Although MME is a comprehensive benchmark, it still has room for improvement in terms of capability coverage, such as covering more scenarios that require reasoning. We will continue to iterate MME series in the future to meet the evaluation requirements of MLLMs. More importantly, we hope that through the design of benchmarks, we can reflect the thoughts on the next capabilities of the model and thereby promote its development.
Acknowledgments
This work is funded by National Natural Science Foundation of China (Grant No. 62506158 and No. 62441234), Fundamental Research Funds for the Central Universities, AI & AI for Science Project of Nanjing University (No. 2024300529), and CCF-Tencent Rhino-Bird Open Research Fund.